# Creative Beam Search: LLM-as-a-Judge for Improving Response Generation

### v2

@@ Introduction
In this paper, we propose a new sampling scheme, namely Creative Beam Search (CBS), to better capture certain key aspects of the creative process. According to the framework proposed in (?), creativity should involve the following steps: task presentation (from internal or external stimuli); preparation; response generation (thanks to creativity-relevant skills); and response validation (thanks to domain-relevant skills). In particular, CBS first simulates the response generation phase through Diverse Beam Search (DBS) (?), generating a more diverse set of possible solutions. Then, it performs a self-evaluation phase in LLM-as-a-Judge style (?) to select the final output. We evaluate our method against the classic sampling strategy with a qualitative assessment study, finding that end-users find our approach preferable and on average CBS provides better solutions than DBS.
@@ Discussion
This paper has introduced a new sampling scheme, Creative Beam Search, to tackle the misalignment between the human creative process and how generative models produce their outputs. It leverages recent techniques such as Diverse Beam Search and LLM-as-a-Judge to simulate aspects of response generation and validation. However, it does not address other key aspects as outlined by (?), such as task motivation from internal stimuli and the possibility of iteratively adjusting the responses. Moreover, both Diverse Beam Search and LLM-as-a-Judge have limitations. For instance, Diverse Beam Search uses Hamming diversity, which only considers differences at the same time step. This can lead to overly similar sequences due to minor misalignments such as initial spacing. In addition, it is only applicable to sequence generation tasks and is more expensive than classic decoding strategies. As for LLM-as-a-Judge, it is important to remark that LLMs are not conscious or intentional. Therefore, self-evaluation does not reflect any personal belief but merely returns what the model has learned to be more likely. Consequently, our approach is just an artificial simulation of certain aspects of creativity.

### v4

@@ Introduction
In this paper, we introduce Creative Beam Search (CBS), a novel generate-and-test sampling scheme designed to artificially replicate certain aspects of the creative process. According to the framework proposed in (?), creativity should involve the following steps: task presentation (from internal or external stimuli); preparation; response generation (thanks to creativity-relevant skills); and response validation (thanks to domain-relevant skills). In particular, CBS first simulates the response generation phase through Diverse Beam Search (DBS) (?), generating a more diverse set of possible solutions. Then, it performs a self-evaluation phase in LLM-as-a-Judge style (?) to select the final output. We evaluate our method against the classic sampling strategy with a qualitative assessment study, finding that end-users find our approach preferable and, on average, CBS (as a generate-and-test approach with DBS) provides better solutions than DBS alone.
@@ Discussion
This paper has introduced a new sampling scheme, Creative Beam Search, to tackle the misalignment between the human creative process and how generative models produce their outputs. It leverages recent techniques such as Diverse Beam Search and LLM-as-a-Judge to simulate aspects of response generation and validation. However, it does not address other key aspects as outlined by (?), such as task motivation from internal stimuli and the possibility of iteratively adjusting the responses. Moreover, both Diverse Beam Search and LLM-as-a-Judge have limitations. For instance, Diverse Beam Search uses Hamming diversity, which only considers differences at the same time step. This can lead to overly similar sequences due to minor misalignments such as initial spacing. In addition, it is only applicable to sequence generation tasks and is more expensive than classic decoding strategies. As for LLM-as-a-Judge, it is important to remark that LLMs are not conscious or intentional. Therefore, self-evaluation does not reflect any personal belief but merely returns what the model has learned to be more likely. Consequently, our approach can be considered as an artificial simulation of certain aspects of creativity. Finally, there is a need to extend the experimental evaluation, considering the impact of the prompt structure on the overall results.
@@
Acknowledgments
The participation and presentation at ICCC’24 was supported by the ISA Doctoral Prize (ISA DP), offered by Istituto di Studi Avanzati, Alma Mater Studiorum Università di Bologna.